Identify and explain the dynamics in rotational motion. Compare the dynamics in rotational


motion to the equivalent translational motion.

1
Expert's answer
2021-12-20T10:31:21-0500

Kinematics of rotational motion explains the motion of particles or objects that perform motion in circular trajectories.


"\alpha=\frac{\omega_f^2-\omega_i^2}{2\phi}↔ a=\frac{v_f^2-v_i^2}{2d}.\\\space\\
\Delta\omega=\alpha t↔ \Delta v=at.\\\space\\
\phi=\frac 12 \alpha t^2↔ d=\frac12at^2."